OpenStudy (anonymous):

when k(x)=x^2+5 how do i determine k^-1(x)

OpenStudy (anonymous):

take y = x^2+5 then x^2 = y-5 \[x=\sqrt{y-5}\] so \[k ^{-1}(x) = \sqrt{x-5}\]